# OmniPose6D: Towards Short-Term Object Pose Tracking in Dynamic Scenes from Monocular RGB

### v1

@@ I Introduction
We build upon previous track-any-point works [16, 17, 18, 19], aiming to bridge the gap between keypoint tracking and object pose estimation. Specifically, keypoint tracking usually focuses on translational or unidirectional motion, whereas object pose tracking encompasses a broader spectrum of motion patterns. Meanwhile, while track-any-point primarily concentrates on long-term single-point tracking, object pose tracking typically employs pairs of points within a brief timeframe to determine pose. Our proposed pose tracking pipeline leverages recent advances in video segmentation [20, 21, 22], taking a sequence of masked images of the object as input. We then initialize tracks of grid samples from the first mask via an off-the-shelf keypoint tracker [19], followed by an uncertainty-aware keypoint refinement network to further improve and select the good tracks. This refinement process selects the most reliable tracks, leading to more accurate pose calculation through a structure-from-motion framework.
A key challenge for 6-DoF object pose estimation is the difficulty of accurately annotating real data, especially at large scales. Synthetic datasets [16, 23] simplify label generation, but often lack out-of-plane rotation or significant occlusion, which are common in real-world data. We thus create a large-scale synthetic dataset, OmniPose6D, that emphasizes object-centric diversity and realistic motion trajectories. Our photo-realistic renderings, combined with motion pattern variety, help bridge the sim-to-real gap, leading to stronger performance that better generalize to real-world scenarios.
@@
We perform experiments on both a held-out portion of our OmniPose6D dataset and HO3D [15] real-world dataset. We compare against a diverse set of baselines, representing recent state-of-the-art methods from related problem settings, including local feature matching, simultaneous localization and mapping (SLAM), and keypoint tracking (Figure 1), combined with a Structure-from-Motion (SfM) framework for computing pose. While we observe that short-term RGB monocular object pose tracking poses significant challenges for the baseline methods, our approach—employing training on our synthetic dataset and a pose tracking pipeline with uncertainty estimation—enables us to achieve meaningful improvements over these existing techniques.
@@ II-A Object Pose Tracking
Continuously determining the 3D position and orientation of objects relative to the camera remains a pivotal challenge in computer vision and robotics. Initially, pose tracking at the instance level was developed using filtering methods or by minimizing inter-frame pose variations within a least-squares framework [24, 25, 26, 3, 1, 27, 2, 28, 29]. Subsequent approaches at the category level employed CAD models for offline training and online template matching [6, 30], evolving to handle novel, unseen objects through techniques like 2D-3D feature matching and 2D-2D image matching [31, 10, 9, 8, 32]. Recently, methods such as BundleTrack [12] and BundleSDF [13] have revisited the SLAM framework, enabling immediate adaptation to new objects, yet remain reliant on depth information. While RGB-D methods demonstrate significant potential, the capabilities of monocular RGB-only, model-free methods have yet to be fully exploited. A notable attempt, PIZZA, utilizes a transformer architecture to predict pose changes from raw images, although it struggles in less structured environments [14]. Our research aims to extend RGB-only methods, enhancing robustness and adaptability for diverse, dynamic environments. Inspired by the progression in 2D object tracking from short-term to long-term contexts [33], we focus on short-term challenges in object pose tracking, proposing novel solutions to increase robustness and versatility in pose tracking technologies.
@@ II-B Keypoint Tracking
Recent advancements in keypoint tracking have seen significant improvements. Particle Video Revisited employs a sliding window approach to enhance point tracking through occlusions [16], while TAP-Vid establishes a benchmark for tracking physical points across video sequences [17]. Building on these, TAPIR combines the matching strategies of TAP-Vid with occlusion handling techniques from PIPs to manage occlusions over extended periods [18]. Diverging from these point-specific methods, CoTracker introduces a collaborative tracking framework using dense optical flow and a transformer network to track point groups effectively, especially useful for keypoints associated with an object from the initial frame mask [19]. We have adapted CoTracker’s approach to refine our keypoint tracking processes, differentiating it from general object pose tracking goals. Moreover, advancements in video segmentation now minimize the need for extensive point visibility analysis, shifting our focus towards precision in critical point tracking for accurate pose estimation [22]. We enhance our model with uncertainty estimation to select well-tracked points, thereby improving the overall accuracy of pose tracking.
@@ II-C Datasets
Keypoint tracking datasets, primarily synthetic, are favored due to the ease of generating precise ground truth labels, such as dense optical flow and point tracking, in controlled environments. FlyingThings introduces objects with randomized poses and velocities and simulates their movement via physics engines, although it suffers from a lack of photorealism, leading to a simulation-to-reality gap [34, 16]. TAP-Vid-Kubric, however, employs advanced Blender rendering techniques to enhance photorealism, using a similar approach as FlyingThings [23]. Conversely, the MegaDepth dataset provides a real-world setting and, though it doesn’t directly offer keypoint pairs, allows their derivation from depth data and camera poses [35]. Despite these resources, there remains a lack in datasets suited for tracking object poses dynamically, particularly in scenarios involving interactions and various rotations. To address this, we introduce OmniPose6D, a synthetic dataset developed for training and evaluating dynamic object pose tracking in interactive settings. This dataset is designed to support future research and promote further developments in dynamic object pose tracking.
@@ III OmniPose6D
Simulator Choice: The photorealism of a simulator’s renderings is critical for reducing the simulation-to-reality (sim2real) gap. Although many modern simulators [40, 41, 23, 42, 39] offer advanced rendering capabilities, we choose BlenderProc [39] due to its user-friendly Python scripting for scene generation and seamless support for BOP-format [43] output–a popular format in the object pose community. This ensures that the annotations we generate–including RGB, depth, object pose, normals, and segmentation–are readily usable for future work.
Mesh Choice: Traditionally, ShapeNet [44] or YCB [45] objects have been common for data generation. ShapeNet [44] is known for its scale and variety, but lacks photorealism and primarily consists of rigid, man-made artifacts. In contrast, YCB objects are manually scanned and of higher quality, but of limited scale. Recent advancements in mesh datasets, like OmniObject3D [38], feature well-scanned meshes in both scale and diversity. We’ve optimized these by simplifying the original meshes with Blender [46] to reduce size for more efficient rendering and also eliminating any invalid meshes. As a result, our dataset generation utilizes over 4,000 meshes across more than 200 categories.
@@ IV Uncertainty-aware Keypoint Refinement
Figure 4: Proposed pipeline. Given a segmentation mask in the first frame, the object is tracked and cropped from a sequence of monocular RGB images with video segmentation [22]. We perform dense grid sampling on the initial mask to provide queries for CoTracker [19] to generate the initial keypoint tracks. An uncertainty-aware keypoint-refinement transformer network refines and selects point tracks, which are used to compute 6-DoF poses within a Structure-from-Motion framework. (We show a sample case from the HO3D dataset [15].)
@@ IV-B Training Data Generation
To enhance frame rate diversity, we initially sample $N$ frames sequentially from the original sequence, ensuring that each interval does not exceed a maximum distance of $K$ frames; we empirically set $N=8$ and $K=4$ in our experiments. This ensures variability in the training data. Object poses are intriguing during hand interaction, often resulting in substantial occlusions. Therefore, we simulate occlusions along the mask boundaries in selected images to mimic these conditions. To ensure adequate useful keypoints in the processed inputs, we employ a co-visibility approach inspired by [47]. For establishing keypoint correspondences, the first view is designated as the reference. We project grid-level points sampled on the reference view’s mask to other views. The projection is deduced from the provided depth maps and pose labels. We also compute the visibility of each keypoint by using the projected depth error and cycle projection error mentioned in [47]. Finally, the input is enhanced with several image augmentation operations, including color jitter, Gaussian blur, motion blur, additive shade, and random flipping. We apply rough segmentation to identify the small region of interest within the image and crop it accordingly. The cropping parameters, centered on the segmented area, are adjusted with random shifts and scaling to ensure robustness. This process standardizes the input and the affine transformation matrix is recorded to map the crop back to the original image context.
@@ IV-C Uncertainty Estimation
While CoTracker [19] emphasizes long-term tracking of individual points, we focus on tracking within shorter time frames. We prioritize the quality of points over their quantity, due to the sensitivity of pose computation to outliers. We incorporate an uncertainty estimation branch designed to predict the level of uncertainty for each keypoint. Instead of directly regressing the L2 distance between the predicted keypoint and its ground truth location in pixel space, which can introduce instability during training, we reformulate it as a classification task with $n_{L}$ predefined levels. We empirically set $n_{L}=5$ with levels $L=[1,3,5,10,\infty]$, adjusting thresholds based on the input image size ratio between inference and training to ensure consistent performance across resolutions. This simple yet effective modification can be seamlessly integrated into the existing network architecture for enhanced performance. As many of the tracking points are often high quality, we utilize a balanced cross-entropy loss for identifying and eliminating outliers.
@@ IV-D Object Pose Estimation
Figure 5: Qualitative comparison of the methods on two datasets: (a) OmniPose6D dataset and (b) HO3D [15] dataset. In both images, we visualize 6-DoF pose tracking results, where the contour (white) is rendered with the estimated pose. Our method tracks the object’s pose successfully in both cases, while other methods perform differently.
@@ V-A OmniPose6D Synthetic Dataset
Setting. OmniPose6D has 21 distinct object categories for evaluation, each featuring a varying number of instances and sequences. To account for the long-tailed distribution, we randomly select 5 sequences from each category, with each sequence comprising 100 frames. In these sequences, an object moves randomly in 3D space relative to the camera against a blank background. Since the original dataset lacks inter-object occlusion, we simulate occlusion randomly in each view (see Section IV-B). This process synthesizes potential occlusions caused by hands or other objects and includes imperfections in segmentation inputs. These images with corrupted masks are then used as inputs in our pipeline. As our task concentrates on short-term window segments, we segment long sequences into shorter clips. To ensure each window has sufficient motion, we use a keypoint tracking method [19] to track the average displacement of randomly selected points, choosing keyframes where the motion surpasses a set threshold. The resulting set of keyframes are then evenly divided into multiple subsets. For our experiments, we set each window to encompass 8 frames.
Results. Table I (top) shows quantitative results, and Figure 8 (a) shows qualitative comparisons. Note we present the results using densely sampled visible Ground Truth (GT) keypoint correspondences, which is more accurate but still imperfect, due to object occlusions and errors introduced by the pose computation process (e.g. coplanar points provide less geometric information). Our method outperforms the baselines by a large margin on all the metrics. LoFTR [49], a local feature-matching method, demonstrates good performance despite not incorporating temporal information. This could be beneficial as the randomly simulated occlusions in each frame make it difficult to use temporal information effectively. DROID-SLAM [50] has more trouble with the test cases, as it focuses on efficiency more than accuracy, and the short-window setting does not favor its ability in large-scale scenes. While TAPIR [18] and CoTracker [19] are both keypoint tracking methods, CoTracker outperforms with its group point tracking, especially in tasks where points are correlated, such as object pose tracking. Our proposed approach builds upon CoTracker for even better results. We attribute this to our comprehensive training dataset, which covers more complex scenes with more random occlusion for object pose tracking, and focuses on the most valuable keypoints to compute the poses, from uncertainty estimation.
@@ V-B HO3D Real Dataset
Results. Quantitative results are shown in Table I (bottom) and qualitative comparisons are illustrated in Figure 8 (b). HO3D exhibits a higher degree of realism, due to the presence of image-level noise like variations in lighting and environmental factors, which may not be fully replicated in a synthetic dataset. The occlusion patterns from the hand are also more natural and consistent compared to the higher occlusion pattern randomness of the synthetic dataset. Echoing Section V-A, our method again achieves the best performance. While the sim2real gap still exists, our large-scale synthetic dataset shows potential to help alleviate the problem. We also observe that a consistent pattern emerges: methods that perform well on synthetic datasets tend to also excel on real-world datasets. This indicates that our custom-generated synthetic dataset serves as an effective proxy for evaluating different methodologies.
VI Conclusion
We have introduced a new task of short-term 6-DoF object pose tracking in dynamic settings from RGB-only input. We focus on a model-free setup, only requiring segmentation of the object in the initial frame. To facilitate the development and evaluation of this task, we created a large-scale synthetic dataset, OmniPose6D, for both training and evaluation. We also developed an uncertainty-aware keypoint refinement network, advancing the accuracy and reliability of object tracking. On both synthetic and real datasets, our approach demonstrates state-of-the-art results compared to existing methods. Future work will focus on improving the pipeline’s speed and exploring more effective methods for keypoint sampling and selection, alongside the simultaneous refinement of keypoints and pose.
APPENDIX
A Model Details
Training. Each training sample involves densely sampling keypoints on the initial mask with a grid size of 3. Due to GPU memory constraints, we randomly select 500 tracks from these samples. Contrary to CoTracker, which assumes stationary track points or fragments for initial track continuation, our approach offers two initial track settings. The first uses the original CoTracker predictions as a starting point for refinement, and the second applies gaussian noise (mean=0, std=1 pixel) to the ground truth keypoint for initial track generation, followed by refinement. This dual-mode initialization allows us to exclusively focus on the refinement task. We also load the pre-trained weights from CoTracker, as shown in Table II. During data pre-processing, we simulate occlusions by drawing perturbed black circles centered on random points along the mask boundary; their radiuses are proportional ([0.1, 0.5]) to the size of the mask.
Following CoTracker [19], the keypoint loss is as follows:
The final loss is determined as follows:
B Baselines
C Benchmark datasets
D Pipeline Details

### v2

@@ I Introduction
We build upon prior track-any-point works [16, 17, 18, 19] to bridge keypoint tracking and object pose estimation. While keypoint tracking focuses on translational or unidirectional motion, pose tracking accounts for broader motion patterns. Track-any-point methods emphasize long-term single-point tracking, whereas pose tracking relies on short-term point pairs for pose estimation. Our pipeline integrates advances in video segmentation [20, 21, 22], using masked object images as input. While segmentation quality may vary, our approach remains robust, as the pose tracking pipeline is driven by keypoint consistency rather than strict segmentation accuracy. We initialize grid sample tracks via an off-the-shelf keypoint tracker [19] and refine them with an uncertainty-aware network, selecting reliable tracks for accurate pose estimation via structure-from-motion.
A major challenge in 6-DoF pose estimation is the difficulty of large-scale real-world annotations. Synthetic datasets [16, 23] ease label generation but often lack realistic out-of-plane rotation and occlusion. To address this, we introduce OmniPose6D, a large-scale synthetic dataset emphasizing object-centric diversity and realistic motion trajectories. Our photo-realistic renderings and varied motion patterns enhance sim-to-real transfer, improving generalization to real-world scenarios.
@@
We perform experiments on both a held-out portion of our OmniPose6D dataset and HO3D [15] real-world dataset. We compare against a diverse set of baselines, representing recent state-of-the-art methods from related problem settings, including local feature matching, simultaneous localization and mapping (SLAM), and keypoint tracking (Figure 1), combined with a Structure-from-Motion (SfM) framework for computing pose. We observe that short-term RGB monocular pose tracking challenges baseline methods, while our synthetic training and uncertainty-aware pipeline achieve meaningful improvements.
@@ II-A Object Pose Tracking
Continuously estimating an object’s 3D position and orientation remains a core challenge in computer vision and robotics. Early instance-level pose tracking relied on filtering or least-squares optimization [24, 25, 26, 3, 1, 27, 2, 28, 29], while category-level methods used CAD models for training and template matching [6, 30], later evolving to feature-based matching for unseen objects [31, 10, 9, 8, 32]. Recent SLAM-inspired approaches [12, 13] enable online adaptation but still rely on depth. While RGB-D methods show promise, monocular RGB-only, model-free techniques remain underexplored. While RGB-based method PIZZA [14] employs a transformer-based framework that processes three consecutive frames for 6D pose estimation, this fixed temporal window can render its predictions sensitive to large inter-frame motions, thereby limiting robustness in highly dynamic environments. In contrast, our approach extends RGB-only methods by concentrating on short-term object pose tracking, which better accommodates abrupt motion changes and complex scene dynamics. Inspired by recent advances in 2D tracking [33], our method integrates adaptive mechanisms designed to enhance reliability and versatility.
@@ II-B Keypoint Tracking
Recent advances in keypoint tracking have improved robustness. Particle Video Revisited enhances tracking through occlusions via a sliding window [16], while TAP-Vid benchmarks physical point tracking [17]. TAPIR refines this by integrating TAP-Vid’s matching with PIPs’ occlusion handling [18]. CoTracker takes a different approach, using dense optical flow and transformers to track point groups [19]. We adapt its framework for keypoint refinement, focusing on precise tracking rather than full object pose estimation. With video segmentation reducing the need for extensive point visibility analysis [22], our method prioritizes selecting well-tracked points through uncertainty estimation, improving pose tracking accuracy.
@@ II-C Datasets
Synthetic keypoint tracking datasets are favored for their precise ground truth labels. FlyingThings simulates object motion with physics engines but lacks photorealism, causing a sim-to-real gap [34, 16]. TAP-Vid-Kubric improves realism using Blender rendering [23], while MegaDepth provides real-world data, deriving keypoints from depth and camera poses [35]. However, existing datasets lack dynamic object pose tracking, especially for interactions and diverse rotations. To address this, we introduce OmniPose6D, a synthetic dataset tailored for training and evaluating pose tracking in interactive settings, fostering future research in dynamic object pose tracking.
Figure 4: Proposed pipeline. Given a segmentation mask in the first frame, the object is tracked and cropped from a sequence of monocular RGB images with video segmentation [22] (Example images are from HO3D [15]). Rather than relying on precise segmentation, we perform dense grid sampling on the first mask to initialize keypoint queries for CoTracker [19], enabling robust tracking of consistent keypoints. An uncertainty-aware transformer further refines and selects reliable tracks, ensuring stable pose estimation even with imperfect segmentation. The selected keypoints are then used to compute 6-DoF poses within a Structure-from-Motion framework.
@@ III OmniPose6D
Simulator Choice: We use BlenderProc [37] for its photorealism, user-friendly Python scripting, and support for BOP-format output, ensuring our generated annotations (RGB, depth, pose, normals, segmentation) are readily usable.
Mesh Choice: Instead of using ShapeNet [40] or YCB [41] alone, we leverage well-scanned, diverse meshes from OmniObject3D [36]. We optimize these meshes with Blender [42] to reduce size and remove invalid ones, resulting in over 4,000 meshes across more than 200 categories.
@@
(a) OmniPose6D dataset
(b) HO3D dataset
Figure 5: Qualitative comparison of the methods on two datasets: (a) OmniPose6D dataset and (b) HO3D [15] dataset. In both images, we visualize 6-DoF pose tracking results, where the contour (white) is rendered with the estimated pose. Our method tracks the object’s pose successfully in both cases, while other methods perform differently. We also demonstrate that our proposed pipeline remains effective even with imprecise segmentation (See the details of SegMask prediction of HO3D).
@@ IV-B Training Data Generation
To enhance frame rate diversity, we initially sample $N$ frames sequentially from the original sequence, ensuring that each interval does not exceed a maximum distance of $K$ frames; we empirically set $N=8$ and $K=4$ in our experiments. This ensures variability in the training data. Object poses are intriguing during hand interaction, often resulting in substantial occlusions. Therefore, we simulate occlusions along the mask boundaries in selected images to mimic these conditions. Specifically, We draw perturbed black circles centered on random points along the mask boundary; their radiuses are proportional ([0.1, 0.5]) to the size of the mask. To ensure adequate useful keypoints in the processed inputs, we employ a co-visibility approach inspired by [43]. For establishing keypoint correspondences, the first view is designated as the reference. We project grid-level points sampled on the reference view’s mask to other views. The projection is deduced from the provided depth maps and pose labels. We also compute the visibility of each keypoint by using the projected depth error and cycle projection error mentioned in [43]. Finally, the input is enhanced with several image augmentation operations, including color jitter, Gaussian blur, motion blur, additive shade, and random flipping. We apply rough segmentation to identify the small region of interest within the image and crop it accordingly. The cropping parameters, centered on the segmented area, are adjusted with random shifts and scaling to ensure robustness. This process standardizes the input and the affine transformation matrix is recorded to map the crop back to the original image context.
@@ IV-C Uncertainty Estimation
To estimate keypoint uncertainty in a probabilistic framework, we reformulate the prediction of the keypoint error as a classification problem with $n_{L}$ predefined levels. For each keypoint $i$, let the true error $e_{i}$ be the Euclidean (L2) distance between its predicted location and the ground truth. We discretize the range of possible errors into $n_{L}$ intervals defined by thresholds
$L=[l_{1},l_{2},\dots,l_{n_{L}}],$
where each $l_{k}$ represents the $k$-th error boundary. We empirically set $n_{L}=5$ with the thresholds
$L=[1,3,5,10,\infty],$
which are adjusted based on the ratio between the input image size at inference and the training image size to ensure consistent performance across resolutions. The ground truth label $y_{i}$ for keypoint $i$ is determined as follows ($l_{0}=0$):
$y_{i}=\begin{cases}k,&\text{if }l_{k-1}\leq e_{i}<l_{k}\quad\text{for }k=1,\dots,n_{L}-1,\\ n_{L},&\text{if }e_{i}\geq l_{n_{L}},\end{cases}$
Our network outputs a set of logits $\{z_{i,k}\}_{k=1}^{n_{L}}$ for each keypoint $i$, which are converted into probabilities via the softmax function:
$p_{i,k}=\frac{\exp(z_{i,k})}{\sum_{j=1}^{n_{L}}\exp(z_{i,j})},\quad k=1,\dots,n_{L}.$
Here, $p_{i,k}$ represents the probability that keypoint $i$’s error falls into the $k$-th error interval.
To train the network, we use a balanced cross-entropy loss between the predicted distribution and the ground truth labels:
$\mathcal{L}_{\text{uncertainty}}=-\sum_{i}\sum_{k=1}^{n_{L}}\mathbb{I}(y_{i}=k)\log p_{i,k},$
where $\mathbb{I}(\cdot)$ is the indicator function.
This formulation allows the network to learn a discrete probability distribution over the error levels, thus providing a probabilistic interpretation of the keypoint uncertainty. The resulting uncertainty estimates are then used to rank and select keypoints, ensuring that only the most reliable ones contribute to the pose estimation process.
@@ V-A OmniPose6D Synthetic Dataset
Setting. OmniPose6D has 21 distinct object categories for evaluation, each featuring a varying number of instances and sequences. To account for the long-tailed distribution, we randomly select 5 sequences from each category, with each sequence comprising 100 frames. In these sequences, an object moves randomly in 3D space relative to the camera against a blank background. Since the original dataset lacks inter-object occlusion, we simulate occlusion randomly in each view (see Section IV-B). This process synthesizes potential occlusions caused by hands or other objects and includes imperfections in segmentation inputs. These images with corrupted masks are then used as inputs in our pipeline. As our task concentrates on short-term window segments, we segment long sequences into shorter clips. To ensure each window has sufficient motion, we use a keypoint tracking method [19] to track the average displacement of randomly selected points, choosing keyframes where the motion surpasses a set threshold. The resulting set of keyframes are then evenly divided into multiple subsets. For our experiments, we set each window to encompass 8 frames. This choice is motivated by the need to balance accuracy and computational efficiency when processing keypoint tracking and bundle adjustment.
Results. Table I (top) shows quantitative results, and Figure 8 (a) shows qualitative comparisons. Note we present the results using densely sampled visible Ground Truth (GT) keypoint correspondences, which is more accurate but still imperfect, due to object occlusions and errors introduced by the pose computation process (e.g. coplanar points provide less geometric information). Our method outperforms the baselines by a large margin on all the metrics. LoFTR [45] excels in per-frame translation and rotation metrics due to its precise local feature matching, which is robust to random occlusions in individual frames. However, its lack of temporal consistency leads to accumulated errors over time, affecting ATE and RPE. DROID-SLAM [46] has more trouble with the test cases, as it focuses on efficiency more than accuracy, and the short-window setting does not favor its ability in large-scale scenes. While TAPIR [18] and CoTracker [19] are both keypoint tracking methods, CoTracker outperforms with its group point tracking, especially in tasks where points are correlated, such as object pose tracking. Our proposed approach builds upon CoTracker for even better results. We attribute this to our comprehensive training dataset, which covers more complex scenes with more random occlusion for object pose tracking, and focuses on the most valuable keypoints to compute the poses, from uncertainty estimation.
@@ V-B HO3D Real Dataset
Results. Quantitative results are shown in Table I (bottom) and qualitative comparisons are illustrated in Figure 8 (b). HO3D exhibits a higher degree of realism, due to the presence of image-level noise like variations in lighting and environmental factors, which may not be fully replicated in a synthetic dataset. The occlusion patterns from the hand are also different, which is more consistent compared to the higher occlusion pattern randomness of the synthetic dataset. Echoing Section V-A, our method again achieves the best performance. While the sim2real gap still exists, our large-scale synthetic dataset shows potential to help alleviate the problem. We also observe that a consistent pattern emerges: methods that perform well on synthetic datasets tend to also excel on real-world datasets. This indicates that our custom-generated synthetic dataset serves as an effective proxy for evaluating different methodologies.
V-C Ablation Studies
We ablate on HO3D [15] for its more realistic settings.
Training Dataset Sources. We examine the impact of different training dataset sources on the performance of our system. We primarily focus on the comparison with CoTracker, which achieves the best overall performance across different test datasets. To mitigate the influence of the uncertainty branch, we substitute the refinement network in Figure 4 with the original CoTracker’s [19] network but vary the training sources. These include: 1) utilizing the original weights, which were trained on the TAP-Vid-Kubric dataset [17]; 2) training the network from scratch using our dataset; 3) using pre-trained weights and then fine-tuning them on OmniPose6D. The results, as presented in Table II, reveal that the combination of using pre-trained weights followed by fine-tuning on our dataset yields superior performance. Furthermore, training from scratch on our dataset also outperforms the approach of using the original weights. This indicates that OmniPose6D is both comprehensive and aptly suited for the task of object pose-tracking. Simultaneously, the TAP-Vid-Kubric [17] dataset effectively captures translational motion. Therefore, we recommend the combined use of both datasets to enhance coverage across wider scenarios.
VI Conclusion
We have introduced a new task of short-term 6-DoF object pose tracking in dynamic settings from RGB-only input. We focus on a model-free setup, only requiring segmentation of the object in the initial frame. To facilitate the development and evaluation of this task, we created a large-scale synthetic dataset, OmniPose6D, for both training and evaluation. We also developed an uncertainty-aware keypoint refinement network, advancing the accuracy and reliability of object tracking. On both synthetic and real datasets, our approach demonstrates state-of-the-art results compared to existing methods. However, our pipeline still builds upon existing successes in segmentation, keypoint tracking, and bundle adjustment modules. Future work will focus on the investigation of a broader range of temporal window settings, exploring the usage of the proposed dataset on more baselines, improving the pipeline’s speed and exploring more effective methods for keypoint sampling and selection, alongside the simultaneous refinement of keypoints and pose.
APPENDIX
A Model Details
Training. Each training sample involves densely sampling keypoints on the initial mask with a grid size of 3. Due to GPU memory constraints, we randomly select 500 tracks from these samples. Contrary to CoTracker, which assumes stationary track points or fragments for initial track continuation, our approach offers two initial track settings. The first uses the original CoTracker predictions as a starting point for refinement, and the second applies gaussian noise (mean=0, std=1 pixel) to the ground truth keypoint for initial track generation, followed by refinement. This dual-mode initialization allows us to exclusively focus on the refinement task. We also load the pre-trained weights from CoTracker, as shown in Table II.
Following CoTracker [19], the keypoint loss is as follows:
The final loss is determined as follows:
B Baselines
C Benchmark datasets
D Pipeline Details